Question One: In a competitive market, the market demand is Qd = 48 - 5P and the market supply is Qs = 7P. A price ceiling of $5 will result in a shortage of

a.

36 units.

b.

24 units

c.

16 units

d.

None of the responses are correct

Question Two: If the government decreased the restrictions on allowable pollution emission levels, then the supply curve for a chemical manufacturer that emits some air pollutants will

a.
become flatter
b.

be become steeper

c.

shift to the right

d.

shift to the left

Question Three: Which of the following is most likely to shift the supply curve for electricity to the right?

a.

Consumers becoming more energy conscious

b.

An increase in income

c.

A decrease in the price of coal, an input to producing electricity

d.

An decrease in the price of natural gas, a substitute source of energy

Question Four: Holding all else constant, as additional firms enter an industry

a.

more output is available at each given price

b.

less output is available at each given price

c.
the same output is available at each given price
d.

Unable to tell

Which of the following is most likely to shift the supply curve for electricity to the right?

a.

Consumers becoming more energy conscious

b.

An increase in income

c.

A decrease in the price of coal, an input to producing electricity

d.

An decrease in the price of natural gas, a substitute source of energy
Answer #1

1) Market price is 48 – 5p = 7p or p = 48/12 = $4. Hence price ceiling of $5 has no effect as market price is lower than ceiling price. Option D

2) It should shift rightwards because the cost of controlling pollution is reduced. Option C

3) A decrease in price of an input reduces overall cost of production. This raises production. Option C

4) More output is produced at the same prices so supply shifts right. Option A

5) Option C